OpenStudy (anonymous):

Justin runs 5 miles in 38 minutes. At the same rate, how many miles would he run in 57 minutes?

OpenStudy (texaschic101):

set up a proportion 5/38 = x/57 (5 miles to 38 min = x miles to 57 minutes) cross multiply 38x = (5)(57) 38x = 285 x = 285/38 x = 7.5 Justin can run 7.5 miles in 57 minutes